Ratios are mostly calculated using data drawn from the financial statements of a firm. However, another group of ratios, called market-based ratios, relate to a firm’s observable market value, stock prices, and book values, integrating information from both the market and the firm’s financial statements.

Consider the case of Blue Hamster Manufacturing Inc.:

Blue Hamster Manufacturing Inc. just reported earnings after tax (also called net income) of $9,750,000, and a current stock price of $12.00 per share. The company is forecasting an increase of 25% for its after-tax income next year, but it also expects it will have to issue 2,900,000 new shares of stock (raising its shares outstanding from 5,500,000 to 8,400,000).

If Blue Hamster’s forecast turns out to be correct and its price-to-earnings (P/E) ratio does not change, what does the company’s management expect its stock price to be one year from now? (Note: Round intermediate calculations to four decimal places. Round the expected stock price to two decimal places.)

$9.82 per share

$12 per share

$7.37 per share

$12.28 per share

One year later, Blue Hamster’s shares are trading at $55.80 per share, and the company reports the value of its total common equity as $16,497,600. Given this information, Blue Hamster’s market-to-book (M/B) ratio is   . (Note: Do not round intermediate calculations.)

Is it possible for a company to exhibit a negative EPS and thus a negative P/E ratio?

Yes

No

Which of the following statements is true about market value ratios?

Low P/E ratios could mean that the company has a great deal of uncertainty in its future earnings.

High P/E ratios could mean that the company has a great deal of uncertainty in its future earnings.
Answer #1

Answer a.

This Year:

Earnings per share = Earnings after tax / Number of shares
Earnings per share = $9,750,000 / 5,500,000
Earnings per share = $1.77

P/E ratio = Current price per share / Earnings per share
P/E ratio = $12.00 / $1.77
P/E ratio = 6.7797

Next Year:

Earnings after tax = $9,750,000 * 1.25
Earnings after tax = $12,187,500

Earnings per share = Earnings after tax / Number of shares
Earnings per share = $12,187,500 / 8,400,000
Earnings per share = $1.45

P/E ratio = Current price per share / Earnings per share
6.7797 = Current price per share / $1.45
Current price per share = $9.82

Answer b.

Book value per share = Total common equity / Number of shares
Book value per share = $16,497,600 / 8,400,000
Book value per share = $1.964

Market-to-book ratio = Market price per share / Book value per share
Market-to-book ratio = $55.80 / $1.964
Market-to-book ratio = 28.41

Answer c.

Yes, it is possible to have a negative EPS, if the company earned a net loss during the year.

Answer d.

Low P/E ratio could mean that the company has a great deal of uncertainty in its future earnings